OpenStudy (anonymous):

any short way to solve this question. integrate (1+cos^2x)sin x dx

OpenStudy (anonymous):

\[\int\limits_{}^{} (1+\cos ^{2}x) \sin x dx\]

OpenStudy (mimi_x3):

Let u=cosx => du/-sinx might work..

OpenStudy (mimi_x3):

\[\int\limits(1+u^{2})sinx*\frac{du}{-sinx} \]

OpenStudy (mimi_x3):

Can you do it from here?

OpenStudy (anonymous):

yes get it